OpenStudy (anonymous):

3. Light coming from a fish makes an incidence angle of 30° to normal under the water. The index of refraction of water is 1.33. A fisherman is looking at the fish through air. At what angle with the normal (θr) will the fish appear to the fisherman? Use Snell’s law: nisinθi = nrsinθr.

OpenStudy (anonymous):

@funscience

OpenStudy (matt101):

HINT: The index of refraction for air is 1.00 (memorize this). Now solving this problem with Snell's Law should be easy!

OpenStudy (anonymous):

yes, you can do it. Just plug in the given information.

OpenStudy (anonymous):

recall that r is for refracted angle and i is for the incident angle.

OpenStudy (anonymous):

i know ni is 1.33 and is nr 1? and i think 0i is 30 degrees

OpenStudy (matt101):

Can you post your calculations?

OpenStudy (anonymous):

well i just figured that from the numbers given in the equation

OpenStudy (matt101):

Look at this: \[1.33 \sin30 =1\sin \theta \]\[\sin \theta=1.33\sin30\]\[\theta = \sin^{-1} (1.33\sin30)\]\[\theta = \space ?\]

OpenStudy (anonymous):

you should always draw a diagram. it really helps

OpenStudy (anonymous):

I can plug that into a scientific calculator to get the answer

OpenStudy (anonymous):

I dont really know how to draw pictures for these types of problems

OpenStudy (anonymous):

i got 0.01215

OpenStudy (anonymous):

look at this diagram

OpenStudy (anonymous):

make sure you're in degrees and not radians

OpenStudy (anonymous):

do you know of a good website for a scietific calculator

OpenStudy (anonymous):

you can download a scientific calculator on your phone. look at this file

OpenStudy (anonymous):

or http://web2.0calc.com/

OpenStudy (anonymous):

oh okay those make it a little more clear

OpenStudy (anonymous):

Im trying to get the answer now

OpenStudy (anonymous):

I cant do it for some reason

OpenStudy (anonymous):

you'll need to just do the calculations. @matt101 already gave you the derivation

OpenStudy (anonymous):

so in the calculator i push sin the (1.33 sin) (30)?

OpenStudy (anonymous):

no, you'll need to get the arcsin of 1.33sin30. so one the webcalc you'll have to press 2nd then asin

OpenStudy (anonymous):

When i put it all in i got 30

OpenStudy (anonymous):

i got 22 degrees

OpenStudy (anonymous):

get yourself a calculator

OpenStudy (anonymous):

that's all :D

OpenStudy (anonymous):

I will lol thanks again :)